OpenStudy (anonymous):

The set of points (1, 3), (3, 5), (4, 7) is collinear true or false?

OpenStudy (anonymous):

Find the slope of the distance between each of the two points using the following formula. \[m=\frac{y _{2}-y_{1}}{x_{2}-x_{1}}\] If the slopes are equal, the points are collinear.

OpenStudy (anonymous):

If I knew how to do that I wouldn't of posted the question..

OpenStudy (anonymous):

Okay, the numbers just mean corresponding coordinates. x2 and y2 should be from the same pair, and x1 and y1 should be from the same pair. For example, using the first pair: m = (5 - 3)/(3-1) m = 2/2 m = 1 Now, find the slope between the other two pairs (pair 2: (3,5) and (4,7), pair 3: (1,3) and (4,7). If both are equal to 1, then they are collinear.